I need help in these two problems:
Problem#1
A group of 3 students is to be selected from a group of 12 students to take part in a class in cell biology.
a) In how many ways can this be done?
b) In how many ways can the group who will not take part be chosen?

Problem #2
If a baseball coach has 5 good hitters and 4 poor hitters on the bench and chooses 3 players at random, in how many ways can he choose at least 2 hiters?

Answer by stanbon(48568)

need help in these two problems:
Problem#1
A group of 3 students is to be selected from a group of 12 students to take part in a class in cell biology.
a) In how many ways can this be done?
Answer: 12C3 = [12*11*10]/[1*2*3] = 220
-----------------------
b) In how many ways can the group who will not take part be chosen?
Answer: 12C9 = 220
------------------------
Problem #2
If a baseball coach has 5 good hitters and 4 poor hitters on the bench and chooses 3 players at random, in how many ways can he choose at least 2 hitters?
Answer: They are all hitters; 5 are better at it than the others.
# of ways to choose at least 2 hitters = 9C2 + 9C3 = 36 + 84 = 120
=============
Cheers,
Stan H.